An electron gun emits 2.0 × 10^16 electrons per second. What electric current does this correspond to?

Grade:11

1 Answers

Deepak Patra
9 years ago
Sol. No. of electrons per second = 2 * 10^16 electrons / sec. Charge passing per second = 2 * 10^16 * 1.6 * 10^-9 coulomb/sec = 3.2 * 10^-9 Coulomb/sec Current = 3.2 * 10^-3 A.